### v1

# Yesnt: Are Diffusion Relighting Models Ready for Capture Stage Compositing? A Hybrid Alternative to Bridge the Gap

Abstract
Volumetric video relighting is essential for bringing captured performances into virtual worlds, but current approaches struggle to deliver temporally stable, production-ready results. Diffusion-based intrinsic decomposition methods show promise for single frames, yet suffer from stochastic noise and instability when extended to sequences, while video diffusion models remain constrained by memory and scale.
We propose a hybrid relighting framework that combines diffusion-derived material priors with temporal regularization and physically motivated rendering. Our method aggregates multiple stochastic estimates of per-frame material properties into temporally consistent shading components, using optical-flow-guided regularization. For indirect effects such as shadows and reflections, we extract a mesh proxy from Gaussian Opacity Fields and render it within a standard graphics pipeline. Experiments on real and synthetic captures show that this hybrid strategy achieves substantially more stable relighting across sequences than diffusion-only baselines, while scaling beyond the clip lengths feasible for video diffusion. These results indicate that hybrid approaches, which balance learned priors with physically grounded constraints, are a practical step toward production-ready volumetric video relighting.
Project Page: https://ejuet.github.io/yesnt/
Figure 1: We integrate diffusion-based decomposition priors with variational methods and image-based lighting to achieve physically plausible, temporally stable relighting of captured volumetric content and seamless compositing into diverse virtual environments
1 Introduction
Volumetric capture stages provide a compelling platform for recording human performances in 3D, enabling applications such as immersive telepresence, virtual production, and mixed-reality content creation. These systems allow for captured performances to be replayed from arbitrary viewpoints and inserted into novel environments. Realizing this potential, however, requires addressing two fundamental challenges: novel view synthesis, to render unseen perspectives from sparse camera inputs, and relighting, to adapt the captured content to new lighting conditions so that it blends naturally into target scenes.
The first challenge, novel view synthesis (NVS), has witnessed extraordinary progress in recent years. Breakthroughs such as Neural Radiance Fields (NeRF) [52] and 3D Gaussian Splatting [36], together with a large body of follow-up work [17], have demonstrated the ability to generate photorealistic renderings from sparse views with remarkable fidelity. By combining multi-view cues with expressive scene representations that capture geometry and appearance, these approaches have transitioned from academic prototypes to production-ready tools, establishing a new foundation for volumetric video pipelines [69, 1].
The second challenge, relighting, has been extensively studied, but remains difficult in capture stages where inputs are sparse, lighting is uniform, and material or illumination cues are limited. While one-light-at-a-time (OLAT) capture can enable high-quality relighting [73, 66], it remains impractical for most applications. Inverse rendering provides a principled formulation, yet becomes ill-posed under limited observations needed for dynamic acquisition. Recent data-driven approaches, including diffusion-based decomposition models, offer a promising alternative by directly predicting material and illumination properties from images [58, 81, 34, 43].
This motivates our central question: Are diffusion relighting models ready for capture stage compositing? While their progress is promising, it remains unclear whether these models can meet the demands of real-world use when integrated into volumetric capture pipelines. In this work, we revisit the relighting problem from a practical perspective with an emphasis on stability. We propose a simple yet effective approach that leverages priors from state-of-the-art decomposition models while avoiding the complexity and instability of full inverse rendering or generative diffusion methods. To this end, we introduce a flow-guided temporal regularization scheme for consistent material estimates and a lightweight screen-space relighting module inspired by real-time graphics pipelines with efficient shadow computation. By combining learned priors with physically informed design choices, our method achieves more stable and accurate relighting, making it better suited for production pipelines in volumetric capture.
Contributions.
In summary, our contributions are
  • An end-to-end relighting pipeline for volumetric capture, where we present a practical, production-ready relighting framework that integrates state-of-the-art image decomposition priors with physically-motivated rendering, enabling consistent and photorealistic results across dynamic sequences.
  • Temporal regularization for dynamic relighting that mitigates flickering and appearance instability, preserving both fine details and smooth lighting transitions over time.
  • Comprehensive evaluation of the usability of diffusion-based relighting approaches for in real-world volumetric capture scenarios, highlighting trade-offs in quality, stability, and computational cost.
2 Related Work
Material Estimation and Relighting
Material estimation and relighting aim to recover surface reflectance properties and scene illumination, enabling realistic appearance synthesis under novel lighting conditions. Classical approaches formulate this task as an inverse rendering problem; owing to its inherent ill-posedness, early methods rely on strong, hand-crafted priors [5, 4, 10, 84]. A closely related direction focuses on estimating material appearance parameterized as a spatially varying BRDF (SVBRDF), again leveraging carefully designed priors [49, 2]. The advent of deep learning has relaxed many of these constraints, enabling SVBRDF estimation from a single texture image [18, 67] or from multi-view observations [19]. More recently, deep learning-based inverse rendering frameworks have coupled differentiable rendering with neural networks to jointly recover geometry, reflectance, and illumination from single-view [76, 39, 40, 61, 72], dual-view [7], or multi-view inputs [16, 46].
Recent years have seen significant progress in neural methods for relighting, which aim to synthesize novel illuminations for existing scenes or objects. Neural Radiance Fields (NeRF) [52] have been extended to support relighting by explicitly modeling material and illumination properties, via intrinsic decomposition of scene properties [8, 83, 60, 64], or via a learned scene representation using One-Light-at-a-Time (OLAT) recordings [73]. The introduction of 3D Gaussian Splatting (3DGS) [36] has led to diverse relighting strategies, via intrinsic decomposition [28, 22, 45, 62, 44], or relighting using OLAT data [20, 6, 48]. Some neural relighting methods avoid 3D reconstruction altogether, operating entirely in image space, e.g. by modelling view- and light-dependent transport as a learned function [82] or training a CNN [74] to predict relit results from sparse lighting samples.
More recently, diffusion models with their strong generative expression have been applied to relighting problems. Among these, Ren et al. proposed an image hormization, where they focus on light adaptation based on the background [58], followed by IC-Light [81] addressing illumination harmonization via text prompt guided background generation. Various works discuss relighting via text prompts [38, 34, 80] or light controlled image generation [79]. Neural Gaffer [34] propose relighting by either text prompt or environment map. While these approaches provide good results for single view images, the state-of-the-art Diffusion Renderer [43] unifies inverse and forward rendering within a video diffusion framework. This approach can reduce temporal inconsistencies – which are inherent in single-image settings, since diffusion models tend to produce different outputs for even slightly different inputs – but its applicability is limited by the large memory requirements. These constraints make it less suitable for high-resolution, multi-view capture stages, where consistency across views is essential. We discuss potential strategies to make its use feasible in such scenarios.
Novel-View Synthesis
Novel-view synthesis aims to generate images of a scene from viewpoints not present in the input observations. Early approaches for novel view synthesis primarily relied on rendering from images, and depth-guided warping techniques [23, 31, 12, 13]. Learning-based methods have since been proposed for novel view synthesis using multiplane images [25, 41, 51, 63] or voxel embeddings [33, 50]. More recently, neural radiance fields became the prevalent approach for novel view synthesis [52]. Since then, many variants focusing on more efficient scene representation have been proposed, such as low-rank factorized 4D grid scene representation [14], sparse 3D voxel grids with spherical harmonics [26], additional non-linearity in grid values [35], and directly optimized dense voxel grids [65]. Beyond static scene acceleration, research has expanded NeRFs to dynamic environments [57, 55, 56, 42, 27], enabling relighting by scene decomposition [8, 83, 9, 64]; improving the rendering of glossy surfaces via alternative radiance parameterizations [68], developing face-specific relighting methods [60], while others proposed scene editing without re-training [78]. Real-time rendering has likewise become a key focus [53, 75, 29], and recent work has moved toward alternative representations such as 3D Gaussian Splatting [36], which replaces implicit fields with explicit point-based primitives for real-time rendering. Subsequent work has further improved both rendering speed [32, 3, 24, 30, 54] and quality [70, 47, 71]. Additional advances have extended Gaussian-based methods beyond basic rendering, such as incorporating subsurface scattering [21] and enabling real-time relighting of head avatars [59]. In parallel, researchers have explored Gaussian-based surface reconstruction techniques [77, 15], further broadening their applicability to geometry recovery tasks.
3 Hybrid Relighting Framework
Figure 2: We optimize a Gaussian Opacity Field [77] from multi-view captures to render RGB, depth and normal maps for the novel views and extract a proxy mesh (left). Using a diffusion decomposition model [43] we extract roughness and metallic maps, which we smooth using an optical-flow guided temporal regularization (top). We render the proxy geometry as a shadow caster in the 3d scene (bottom) and blend it with out screen space rendered image (right).
Our method combines three key components as shown in Fig. 2: volumetric scene representations for novel view synthesis and proxy geometry (Sec. 3.1), estimation and regularization of intrinsic material properties (Sec. 3.2 and Sec. 3.3) and a physically based relighting stage that produces view-consistent shading under novel illumination (Sec. 3.4). The final compositing step accounts for indirect interactions such as cast shadows and reflections, ensuring that the subject integrates plausibly with its surroundings. In the following, we describe each of these stages in detail.
3.1 Novel View Synthesis & Indirect Interactions
Given a multi-camera capture of a performance, our goal is to render the subject from novel viewpoints and to integrate it into a new environment with plausible indirect interactions, specifically cast shadows on nearby receivers and reflection contributions, while supporting relighting. To achieve this, we ideally need a NVS representation that provides (i) RGB appearance, (ii) depth maps, (iii) surface normals, and (iv) a proxy geometry suitable for integration into conventional rendering engines. Gaussian Opacity Fields (GOF) [77] are a natural choice, as they jointly optimize for appearance and normals, produce accurate depth estimates, and allow mesh extraction for downstream use.
However, we observe that GOF reconstructions often suffer from artifacts: curved, natural surfaces are approximated as piecewise planar surfaces, and spurious holes may appear. To mitigate these issues, we introduce two additional regularization terms during optimization: a Huber-regularized total variation loss on the rendered depth map, and an edge-aware Laplacian smoothness prior on the Gaussian normal map. Let $z(x,y)$ denote the rendered depth map. We enforce spatial smoothness with a Huber loss–based total variation penalty:
$\mathcal{L}_{\text{depth}}=\sum_{x,y}H_{\delta}(\tfrac{\partial z}{\partial x})+H_{\delta}(\tfrac{\partial z}{\partial y})$ (1)
where $H_{\delta}$ denotes Huber loss (cf. Eq. 6). To prevent blocky surface approximations while respecting appearance discontinuities, we penalize the Laplacian of the estimated normal field $n(x,y)$ in an edge-aware fashion with respect to the RGB image $I(x,y)$ as
$\mathcal{L}_{\text{normal}}=\sum_{x,y}\exp\left(-\tfrac{\|\nabla I(x,y)\|^{2}}{2\sigma^{2}}\right)\cdot\|\Delta n(x,y)\|^{2},$ (2)
where $\sigma$ is a contrast sensitivity parameter. Intuitively, smoothing is encouraged in homogeneous regions while being suppressed at edges aligned with appearance boundaries.
Proxy Mesh for Indirect Interactions.
The extracted GOF mesh enables conventional rendering of indirect effects in external engines such as Blender. By inserting the mesh into the background scene and disabling intersections with foreground camera rays, we obtain plausible cast shadows and reflection contributions without requiring perfect geometry or material accuracy. Since indirect effects are typically low-frequency, coarse geometry suffices for the majority of capture scenarios to render plausible background images.
3.2 Hybrid G-Buffer Estimation
We estimate G-buffer channels using learned predictions for view- and light-independent material properties. We consider the problem of predicting roughness $r$ and metallic $\mu$ maps, from a single color image $I\in\mathbb{R}^{H\times W\times 3}$ by employing the intrinsic decomposition model $\mathcal{D}_{\theta}$ from [43] that provides the mapping
$\mathcal{D}_{\theta}:I\rightarrow\{r,\mu\}.$ (3)
Since diffusion sampling is stochastic, we draw $K$ samples per input in order to estimate both mean predictions and per-pixel uncertainty. To allow high-resolution processing, inference is carried out on overlapping patches $I^{(j)}$ as
$\hat{x}^{(i,j)}=\mathcal{D}_{\theta}(I^{(j)};\,\xi_{i}),\quad i=1,\dots,K$ (4)
where $\xi_{i}$ denotes the diffusion noise. These patchwise predictions are then blended into full-resolution maps $\tilde{x}^{(i)}$, yielding an ensemble $\{\tilde{x}^{(1)},\dots,\tilde{x}^{(K)}\}$ from which we compute per-pixel means and variances, thereby suppressing sampling variability and exposing an uncertainty signal useful for downstream processing.
3.3 Temporal Consistency via Optical Flow-Guided Regularization
We next enforce consistency across neighboring renderings. Let $\tilde{x}\in\{\tilde{r},\tilde{\mu}\}$ denote the per-pixel mean prediction and $\sigma_{x}$ the corresponding normalized standard deviation. Let $m_{\alpha}=\mathbbm{1}_{\alpha>0.5}$ be a binary foreground mask, and let $\mathcal{W}$ denote backward warping from frame $t+1$ to $t$ using estimated optical flow. We write $\odot$ for element-wise multiplication and $\mathcal{V}$ to denote a binary map of valid flows which point to valid foreground $\mathcal{W}(m_{\alpha})$ and are forward/backward consistent. The temporally regularized estimate $x_{t}$ at step $t$ is obtained by minimizing
$\begin{split}\underset{x}{\arg\min}&\ \sum_{t}\langle H_{\delta}(\tilde{x}_{t}-x_{t}),1-\sigma_{x,t}\rangle+\lambda_{1}TV_{\varepsilon}(x_{t})\\ &+\lambda_{2}\left\|\left(\mathcal{W}(x_{t})-x_{t}\right)\odot\mathcal{V}\right\|_{2}^{2}\end{split}$ (5)
where $\langle\cdot,\cdot\rangle$ is the element-wise inner product and $H_{\delta}$ is the (element-wise) Huber loss
$H_{\delta}(a)=\begin{cases}\frac{1}{2}a^{2}&\text{if }|a|<\delta\\ \delta\left(|a|-\frac{1}{2}\delta\right)&\text{otherwise.}\end{cases}$ (6)
The (smoothed) isotropic total variation is
$\mathrm{TV}_{\varepsilon}(x^{t})\;=\;\sum_{i}\sqrt{\|\nabla_{x}x^{t}_{i}\|_{2}^{2}+\|\nabla_{y}x^{t}_{i}\|_{2}^{2}+\varepsilon},$ (7)
for small $\varepsilon$. The first term in Eq. 5 is a confidence-weighted data fidelity loss that aligns $x$ with the diffusion prediction $\tilde{x}$, with deviations weighted according to prediction confidence (i.e., lower weight in regions of high uncertainty). The second term enforces spatial smoothness through total variation regularization, and the third term constrains temporal consistency by penalizing discrepancies between the current roughness or metallic estimate and its backward-warped counterpart.
3.4 Relighting and Compositing
We perform single-image environment relighting using albedo, depth, normals, roughness and metallic estimated as described above. Shading is computed using Disney’s Principled BRDF [11] using a Cook–Torrance microfacet BRDF with a GGX normal distribution, Smith geometry term, and Schlick Fresnel approximation. Illumination is provided by a high-dynamic-range environment map rendered from the target scene at the approximate actor location. For each environment direction, we evaluate the BRDF contribution modulated by a soft shadow term, which is computed by ray-marching from each surface point along the light direction in view space. At each step, the sample position is projected back into screen space, and the depth buffer is queried to detect occluders. An occlusion factor is accumulated over all steps, then normalized to produce a soft shadow map in the range $[0,1]$.
We composite the gamma-corrected, relit foreground with the target background using the alpha channel from the GOF render.
4 Experiments
We analyze the performance of diffusion-based methods for relighting across multiple scenes and lighting conditions, and demonstrate that our approach offers advantages in terms of robustness and visual reliability. Our evaluation comprises qualitative results on both real volumetric recordings and synthetic data (Sec. 4.1), complemented by a quantitative analysis on the synthetic dataset (Sec. 4.2). To further demonstrate the effectiveness of individual components of our method, we include an ablation study (Sec. 4.3), and discuss limitations in Sec. 4.4.
For comparison, we benchmark our approach against several recent state-of-the-art methods in neural relighting and rendering, including NVIDIA Diffusion Renderer [43], Neural Gaffer [34], IC-Light [81], Gaussian Opacity Fields (GOF) [77], and Relightable 3D Gaussian (R3DG) [28].
Datasets
Our real-world evaluation is based on an extended set of various scenes from the RIFTCast dataset [85]. Each scene consists of 36 high-resolution color images at a resolution of $24.5$ MP, captured within a calibrated light stage with cameras arranged in a cylindrical rig of 5.25 meters in diameter. We constructed $12$ diverse target environments in Blender to place the captured content.
[…]
Implementation Details
All experiments are implemented in PyTorch. We use the Adam optimizer [37] with its default parameters.
4.1 Qualitative Evaluation
[…]
To further highlight the improvements over GOF, we provide an additional comparison in Fig. 4, focusing on challenging regions such as hair. Our method recovers clean alpha values instead of relying on coarse mesh representations, resulting in more natural boundaries. We also observe smoother surface reconstruction, which is particularly noticeable in the arm region, where GOF often produces sharp edges.
Finally, we also report qualitative results on the synthetic dataset in Fig. 5. These examples show ground-truth renderings alongside our relit outputs.
GOF Mesh [77]
IC-Light [81]
Neural Gaffer [34]
R3DG [28]
Diff. Renderer [43]
Yesnt (Ours)
Figure 3: Qualitative comparison with recent state-of-the-art relighting methods. Each row corresponds to a different scene under novel lighting conditions. Top panels show the full relit view, while bottom panels provide close-up crops.
Figure 4: Comparison of relighting quality between GOF meshes [77] (left) and our method (right). GOF often approximates natural regions such as the arm with piecewise planar surfaces, whereas our approach produces smoother reconstructions and cleaner alpha mattes in challenging areas like hair.
Figure 5: Qualitative results on the synthetic dataset, showing ground truth (top) and our method (bottom).
4.2 Numerical Analysis of Relighting Accuracy
Table 1: Quantitative comparison of relighting accuracy on synthetic data for diffusion-based methods. GOF is included for reference.
Method Conditioning PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$
GOF Mesh [77] HDR Env. Map $29.4181\pm 5.7897$ $9458\pm 0.0397 0.0$$0619\pm 0.0552$
IC-Light [81] LDR Background $19.4115\pm 4.5775$ $9078\pm 0.0810 0.0$$0873\pm 0.0839$
Neural Gaffer [34] HDR+LDR Env. Map $24.4263\pm 4.5009$ $9166\pm 0.0837 0.0$$0836\pm 0.0918$
Diffusion Renderer [43] LDR+Log Env. Map $27.4926\pm 6.1947$ $9374\pm 0.0501 0.0$$0817\pm 0.0710$
Yesnt (Ours) HDR Env. Map $30.1918\pm 5.4833$ $9580\pm 0.0284 0.0$$0523\pm 0.0462$
We complement our qualitative evaluation with a quantitative analysis on object relighting on the synthetic dataset. To this end, we render ground-truth references under a wide range of environment maps and measure reconstruction accuracy across competing methods. Table 1 summarizes the results across standard image-based metrics.
Our method consistently achieves the best scores across all three metrics. In particular, it improves PSNR by a clear margin over all competing diffusion-based approaches, while also yielding the highest SSIM and lowest LPIPS. GOF serves as a strong geometric baseline, yet it is limited by hard surface boundaries and struggles in regions with complex transparency, as already seen qualitatively. In contrast, our pipeline maintains both high fidelity and perceptual quality across diverse lighting conditions.
4.3 Ablation Study
We conduct an ablation study to assess the contribution of individual components of our model. Tab. 2 reports results when removing either the relighting branch or the estimation of G-Buffer maps (roughness/metallic). Without relighting, the network reduces to a pure NVS model and achieves substantially lower accuracy. Removing the intrinsic decomposition and keeping roughness and metallic constant also results in a notable drop across all metrics, showing that physically meaningful intermediate representations are crucial for high-fidelity relighting. The full model achieves the highest scores across all metrics.
We further evaluate the effect of temporal regularization, as introduced in Sec. 3.3. Tab. 3 compares the full model against a variant without temporal regularization. The latter suffers from degraded temporal PSNR (tPSNR) and a higher perceptual path length (PPL), indicating the presence of flickering artifacts across video frames. Note, that while PPL is typically defined in the context of latent interpolations in generative models, we compute it directly across consecutive output frames to measure frame-to-frame perceptual variation. Fig. 6 illustrates this flickering effect on predicted roughness maps: the Diffusion Renderer [43] exhibits strong temporal inconsistency across consecutive frames, whereas our flow-guided smoothing produces temporally coherent results while preserving spatial detail. These findings underline the importance of temporal regularization for stable relighting in dynamic scenes.
Table 2: Ablation study of our model, comparing variants without any relighting (i.e. novel view synthesis only), without estimation of G-Buffer maps (roughness and metallic), and the full model.
Variant PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$
No Relighting $22.8414$ $0.9265$ $0.0742$
No Decomposition $28.5688$ $0.9507$ $0.0565$
Full $30.1918$ $0.9580$ $0.0523$
Table 3: Temporal regularization ablation on videos (see Sec. 3.3). Removing temporal regularization (w/o Temporal Reg.) leads to a decrease in temporal PSNR (tPSNR) and larger PPL, indicating stronger flickering artifacts.
Variant tPSNR $\uparrow$ PPL $\downarrow$
w/o Temporal Reg. $28.8737$ $0.3598$
Full $29.3660$ $0.3587$
Table 4: Runtimes Analysis of our method compared to others. Bottlenecks are the Gaussian model optimization and the diffusion evaluations.
Step Runtime
Per Scene
GOF – Optimization [77] 1.1 h
R3DG – Optimization [28] 0.9 h
Per Frame
GOF – Rendering 0.4 s
Diffusion Decomposition 80 s
Temporal Regularisation 1.5 s
Relighting 1.5 s
Full Pipeline 87 s
IC-Light [81] 14 s
Neural Gaffer [34] 18 s
R3DG – Relighting [28] 1 s
Diffusion Renderer [43] 28 s
Figure 6: Top: Roughness maps predicted by the Diffusion Renderer [43] across three consecutive frames, showing strong temporal inconsistency. Bottom: Our flow-guided smoothing produces temporally coherent roughness estimates while preserving spatial detail.
4.4 Limitations
In terms of efficiency, the runtime of our method is dominated by the diffusion decomposition stage (cf. Tab. 4). Our method assumes a single environment‐map illumination and does not model spatially varying light sources, which limits local light transport effects (e.g., missing cast shadows on limbs). Finally, our approach relies on diffusion predictions being approximately correct on average. When the model yields inconsistent hypotheses across frames, the temporal regularizer suppresses these variations, leading to over-smoothing and a loss of fine detail and contrast.
5 Conclusion
We asked: Are diffusion relighting models ready for capture stage compositing? While they produce impressive results, they do not scale well to longer sequences. By aggregating stochastic material estimates and enforcing temporal coherence with flow-guided smoothing, our method addresses this shortcoming. Experiments on real and synthetic captures show higher fidelity and reduced flicker, bridging the gap between prototypes and production pipelines. In short, the answer is Yesnt: diffusion models alone are not yet ready, but our hybrid approach offers a practical path toward production-ready volumetric video relighting.
Acknowledgments
This work has been funded by the Ministry of Culture and Science North Rhine-Westphalia under grant number PB22-063A (InVirtuo 4.0: Experimental Research in Virtual Environments). We are deeply grateful to Janelle Pfeifer for her invaluable assistance during data capture and processing.